# LASER: Attention with Exponential Transformation

### v1

@@ Abstract
Transformers have had tremendous impact for several sequence related tasks, largely due to their ability to retrieve from any part of the sequence via softmax based dot-product attention. This mechanism plays a crucial role in Transformer’s performance. We analyze the gradients backpropagated through the softmax operation in the attention mechanism and observe that these gradients can often be small. This poor gradient signal backpropagation can lead to inefficient learning of parameters preceeding the attention operations. To this end, we introduce a new attention mechanism called LASER, which we analytically show to admit a larger gradient signal. We show that LASER Attention can be implemented by making small modifications to existing attention implementations. We conduct experiments on autoregressive large language models (LLMs) with upto 2.2 billion parameters where we show upto 3.38% and an average of $\sim$1improvement over standard attention on downstream evaluations. Using LASER gives the following relative improvements in generalization performance across a variety of tasks (vision, text and speech): 4.67% accuracy in Vision Transformer (ViT) on Imagenet, 2.25% error rate in Conformer on the Librispeech speech-to-text and 0.93% fraction of incorrect predictions in BERT with 2.2 billion parameters.
@@ 1 Introduction
We conduct thorough empirical verification across a variety of Transformer models: Conformer (Gulati et al., 2020) for Librispeech speech-to-text (Panayotov et al., 2015), Vision Transformer(Dosovitskiy et al., 2021) for ImageNet classification (Deng et al., 2009), decoder-only text Transformer (Brown et al., 2020) on C4 dataset (Raffel et al., 2020) and BERT ((Devlin et al., 2018)). We conduct experiments on decoder-only autoregressive language models from 234 million parameters to 2.2 billion parameter models, where we demonstate improvements of up to 1.7% relative improvement in test loss over standard attention. We conduct one-shot evaluation on 17 downstream tasks and show that LASER outperforms standard attention on 14 tasks with upto 3.38% improvement in accuracy and 1% accuracy improvement on average. On a 2.2 billion parameter BERT (Devlin et al., 2018), LASER gives a relative improvement of 0.93% on masked language modeling prediction error rate. LASER also demonstrates a 4.67% relative improvement in validation error rate in Vision Transformer and 1.2% absolute improvement in accuracy, and a 2.25% relative improvement in validation word error rate in the Conformer benchmark.
@@ 2 Related Work
The Mamba architecture introduces state-space models (SSMs) as a replacement for traditional attention. Models like S6 (S4+selection+scan) (Gu & Dao, 2023) from Mamba and SSD from Mamba-2 (Dao & Gu, 2024) have linear computational complexity in sequence length without the use of attention. However, despite these innovations, attention-based models like Gemini 1.5 Flash (Gemini, 2024) and LLaMA 3 (Dubey et al., 2024) continue to dominate long context regime, particularly through advancements in context parallelism (Liu et al., 2023), which ensures scalability while maintaining the strengths of attention mechanisms in Transformer models.
Efficient attention mechanisms have become critical in handling long sequences, especially in Transformer-based architectures. This mechanism is used for faster inference and training, particularly when scaling up to large sequence lengths. Sparse Transformers (Child et al., 2019) use fixed sparse attention patterns, enabling them to efficiently handle very long sequences by reducing the quadratic complexity of standard attention to linear or sub-quadratic in practice. Routing Transformers (Roy et al., 2021) take a different approach by introducing a mechanism that sparsifies attention through data-dependent sparsity patterns with subquadratic computational complexity in sequence length. Similarly, Longformer (Beltagy et al., 2020) modifies the standard self-attention mechanism to handle long documents by combining local windowed attention with global attention. FlashAttention (Dao et al., 2022; 2024) is a recent advancement that reduces HBM memory accesses by using GPU SRAM during attention computation, improving the speed of attention computation. LASER Attention can be thought of as complementing these approaches, as it conducts attention using the exponential transformation of inputs, without any change to the underlying attention function.
@@ 3.1 Transformers and Softmax Dot-Product Attention
Let $X\in\mathbb{R}^{N\times d}$ represent the input sequence with $N$ tokens, where the $i$-th row is a $d$-dimensional representation of the $i$-th token. We describe the Transformer layer $T:\mathbb{R}^{N\times d}\to\mathbb{R}^{N\times d}$ similar to Katharopoulos et al. (2020) as follows:
@@
where the inputs to the network $Z\in\mathbb{R}^{N\times V}$. Thus choosing a suboptimal attention function $\operatorname{attn}(.)$ can affect every layer of the final $\operatorname{Transformer}(.)$. Let $\ell(\operatorname{Transformer}(Z),Y)$ be the loss function used to learn the parameters of the Transformer, where $Y$ is the true label information. Autoregressive language modeling (Radford et al., 2018; Brown et al., 2020) involves using a causal mask $M$, which is a lower triangular matrix, and is multiplied before the softmax operation as follows:
$\operatorname{attn}(X)$ $=\operatorname{softmax}(M\odot QK^{\top})V,$
$M_{ij}$ $=\mathbbm{1}\{i\leq j\}$
@@ 3.2 Gradient Analysis of Attention
For simplicity, we first let the sequence length $N$ be $2$ with attention probabilities $A=\operatorname{softmax}(KQ^{\top})$ and attention logits as $\tilde{A}=KQ^{\top}$. Expanding the matrices $A$ and $\tilde{A}$, we get:
@@
Thus, small Jacobian magnitude can lead to small backpropagated gradient. We now analyze an element of the Jacobian:
attn $\operatorname{attn}\text{ Jacobian:}\ \ \frac{\partial o_{1}}{\partial\tilde{a}_{11}}$ $=v_{1}\sigma(\tilde{a}_{11}-\tilde{a}_{12})(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))-v_{2}\sigma(\tilde{a}_{11}-\tilde{a}_{12})(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))$
@@
$\implies\operatorname{laser}(X)$ $=\log(\operatorname{softmax}(QK^{\top})\exp(V))\quad\rightarrow\quad\text{LASER Attention},$ (6)
where $\log(.)$ and $\exp(.)$ are applied elementwise. Expanding (6) for $N=2$ and $d=1$ as done for standard attention (4) gives:
@@
Low gradient saturation. Computing an element in the Jacobian ${\partial\operatorname{laser}(X)}/{\partial\tilde{A}}$ as done for standard attention in (5) gives the following:
[…]
$=\frac{(\exp(v_{1})-\exp(v_{2}))\sigma(\tilde{a}_{11}-\tilde{a}_{12})(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))}{\sigma(\tilde{a}_{11}-\tilde{a}_{12})(\exp(v_{1})-\exp(v_{2}))+\exp(v_{2})}.$
We now consider a limiting case to simplify the above equation: if $v_{1}-v_{2}\gg 0$, then
[…]
$\approx\underbrace{(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))}_{\text{low saturation}},$
@@
Lemma 3.2 (Boyd & Vandenberghe (2004)).
@@
$\max\{x_{1},\dots,x_{n}\}\leq f(x_{1},\ldots,x_{n})\leq\max\{x_{1},\dots,x_{n}\}+\log n$
@@ 3.3 LASER Implementation via Log-Weighted-Sum-Exp Trick
$o_{1}=\log(\sigma(\tilde{a}_{11}-\tilde{a}_{12})\exp(v_{1})+(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))\exp(v_{2})),$
one can notice that $\exp(.)$ operations can lead to overflow. This problem has been recognized in Blanchard et al. (2019) and the “log-sum-exp trick” is used to avoid overflows. However, the log-sum-exp trick cannot be applied directly as it would be difficult to implement without changing the underlying attention function. We propose a “log-weighted-sum trick”, where we subtract the maximum value $m=\max(v_{1},v_{2})$ from $v_{1}$ and $v_{2}$ and rewrite the above equation as follows:
$o_{1}$ $=\log((\sigma(\tilde{a}_{11}-\tilde{a}_{12})\exp(v_{1}-m)+(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))\exp(v_{2}-m))*\exp(m))$
$=\log(\sigma(\tilde{a}_{11}-\tilde{a}_{12})\exp(v_{1}-m)+(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))\exp(v_{2}-m))+m.$
@@
Define ON×d as: $\text{Define $O\in\mathbb{R}^{N\times d}$ as: }O$ $=\log(\operatorname{softmax}(QK^{\top})\exp(\hat{V})\operatorname{diag}(\exp(m)))$
$O_{ij}$ $=(\log(\operatorname{softmax}(QK^{\top})\exp(\hat{V})))_{ij}+m_{j}.$ (9)
Here, $m=(m_{1},\ldots,m_{d})$ and $\operatorname{diag}(m)$ is a diagonal matrix with elements of $m$ as diagonals. Log-weighted-sum-exp trick allows us to implement LASER attention via merely modifying the inputs and outputs of standard attention, without changing the underlying attention function. Additionally, we show in Section 4.1 that this trick helps avoid overflows in large scale settings. The following JAX (Bradbury et al., 2018) code demonstrates that LASER attention can be implemented by utilizing standard attention functions.
@@
$\hat{V}\in\mathbb{R}^{N\times d}\text{ such that }\hat{V}_{ij}=(V_{ij}-m_{j})\quad\text{// Shift values to avoid overflow in the following }\exp(.)$
@@
Define ON×d as: ($\text{Define $O\in\mathbb{R}^{N\times d}$ as: }(O)_{ij}=(\log(\operatorname{softmax}(QK^{\top})\exp(\hat{V})))_{ij}+m_{j}$
6:  return  LASER attention output $O$
@@ 4.1 Autoregressive Language Modeling on C4
(a) Training Loss Comparison
(b) Test Loss Comparison
Figure 2: Comparison between pretraining performance of LASER and standard attention mechanism on a 301 million parameter autoregressive language model with 32 layers, 2048 hidden dimension and 1024 model dimension on 168 billion tokens in C4 dataset. LASER attention shows consistently lower train (left) and test (right) loss values.
@@
To demonstrate the scalability of our approach, we conducted experiments on a 1.1 billion and 2.2 billion parameter model. We note that without using the log-weighted-sum-exp trick introduced in Section 3.3, the 2.2 billion parameter model training fails due to numerical overflow. In Figure 4, we show that LASER attention outperforms standard-attention in a 2.2 billion parameter model with model dimension 2048 and hidden dimension 8192 with 32 layers and 8 attention heads (each of size 512). We observe the same in 1.1 billion model (Figure 4), which has a scaled down hidden dimension (4096) and attention head size (256).
@@
In Figure 2, we mention the performance of our 2.2 billion parameter model on several downstream tasks. Where we evaluate on ARC (Clark et al., 2018), BoolQ (Clark et al., 2019), CB (Wang et al., 2019), COPA (Wang et al., 2019), HellaSwag (Zellers et al., 2019), MultiRC (Khashabi et al., 2018), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), RACE (Lai et al., 2017), ReCoRD (Zhang et al., 2018), RTE (Wang et al., 2019), StoryCloze (Mostafazadeh et al., 2016), WiC (Pilehvar & Camacho-Collados, 2019), Winograd (Levesque et al., 2012), Winogrande (Kocijan et al., 2020), and WSC (Wang et al., 2019). We found that LASER outperforms in 14 out 17 datasets with upto 3.38% difference and 0.85% difference on average in accuracy.
WSC Winogrande Winograd WiC StoryCloze RTE ReCoRD
LASER 81.40 61.88 81.68 51.41 77.87 53.07 85.39
Standard 79.65 62.27 80.22 50.94 76.59 53.07 85.14
RaceM RaceH PIQA OpenBookQA MultiRC HellaSwag COPA
LASER 50.56 37.71 77.26 48.80 57.51 66.56 82.00
Standard 49.44 37.65 76.88 47.60 54.13 65.49 80.00
CB BoolQ ARC_Easy
LASER 41.07 63.52 61.53
Standard 44.64 60.58 60.82
Table 2: Accuracies of one-shot evaluation of 2.2 billion parameter autoregressive language model trained via LASER and standard attention. We found that LASER outperforms standard attention on 14 out of 17 datasets by upto 3.38%. On average LASER gives an accuracy of 63.48% vs standard attention (62.65%).
(a) Train loss - 2.2 billion parameter model.
(b) Test loss - 2.2 billion parameter model.
(c) Train loss - 1.1 billion parameter model.
(d) Test loss - 1.1 billion parameter model.
Figure 4: Performance comparison for 2.2 billion and 1.1 billion parameter models. The 2.2 billion model has 32 layers, 8 attention heads with head size 512, MLP hidden dimension 8192, and model dimension 2048. The 1.1 billion model has 32 layers, 8 attention heads (head size 256), MLP hidden dimension 4096, and model dimension 1024. LASER demonstrates better train and test loss compared to Standard Attention even in large scale setting.

### v2

@@ Abstract
Transformers have had tremendous impact for several sequence related tasks, largely due to their ability to retrieve from any part of the sequence via softmax based dot-product attention. This mechanism plays a crucial role in Transformer’s performance. We analyze the gradients backpropagated through the softmax operation in the attention mechanism and observe that these gradients can often be small. This poor gradient signal backpropagation can lead to inefficient learning of parameters preceeding the attention operations. To this end, we introduce a new attention mechanism called LASER, which we analytically show to admit a larger gradient signal. We show that LASER attention can be implemented by making small modifications to existing attention implementations. We conduct experiments on autoregressive large language models (LLMs) with upto 7.7 billion parameters with an average improvement of upto 1.44% over standard attention on downstream evaluations and 1.65% finetuning improvements. Additionally, LASER demonstrates generalization performance improvement across a variety of tasks (vision, text and speech):Vision Transformer (ViT) on Imagenet, Conformer on the Librispeech speech-to-text and BERT with 2.2 billion parameters.
Machine Learning, ICML
@@ 1 Introduction
We conduct thorough empirical verification across a variety of Transformer models: Conformer (Gulati et al., 2020) for Librispeech speech-to-text (Panayotov et al., 2015), Vision Transformer(Dosovitskiy et al., 2021) for ImageNet classification (Deng et al., 2009), decoder-only text Transformer (Brown et al., 2020) on C4 dataset (Raffel et al., 2020) and BERT ((Devlin et al., 2018)). We conduct experiments on decoder-only autoregressive language models from 234 million parameters to 7.7 billion parameter models, where we demonstate improvements of up to 1.7% relative improvement in test loss over standard attention. We conduct one-shot evaluation on several downstream tasks and show that LASER outperforms standard attention with upto 1.44% accuracy improvement on average and 1.65% improvement on average upon finetuning. On a 2.2 billion parameter BERT (Devlin et al., 2018), LASER gives a relative improvement of 0.93% on masked language modeling prediction error rate. LASER also demonstrates 1.2% improvement in accuracy, and 0.2% improvement in validation word error rate in the Conformer benchmark.
@@ 2 Related Work
A relevant work (Veličković et al., 2024) studies the importance of expressing sharp distributions in attention mechanisms of transformer-based large language models. This helps with downstream tasks such as text retrieval. In contrast, our paper finds that sharp distributions in softmax lead to gradient bottleneck during backpropagation and proposes LASER to fix this issue. In large vision transformer (Dosovitskiy et al., 2021) pretraining, sharp softmax attention distributions led to training instabilities in (Dehghani et al., 2023), the authors propose QK-normalization, which applies layer normalization to queries and keys before computing attention weights. This equilibrates the norms of all the queries and keys, leading to uniform distributions. In (Kim et al., 2021), the authors identified that standard attention mechanism is not $\ell_{2}$-Lipschitz and develop $\ell_{2}$ multi-head attention, which uses $\ell_{2}$-divergence instead of dot-product between key and query vectors. This formulation admits a Lipschitz constant.
@@ 3.1 Transformers and Softmax Dot-Product Attention
Let $X\in\mathbb{R}^{N\times d}$ represent the input sequence with $N$ tokens, where the $i$-th row is a $d$-dimensional representation of the $i$-th token. We describe the Transformer layer $T:\mathbb{R}^{N\times d}\to\mathbb{R}^{N\times d}$ similar to (Katharopoulos et al., 2020) as follows:
@@
where the inputs to the network $Z\in\mathbb{R}^{N\times V}$. Thus choosing a suboptimal attention function $\operatorname{attn}(.)$ can affect every layer of the final $\operatorname{Transformer}(.)$. Let $\ell(\operatorname{Transformer}(Z),Y)$ be the loss function used to learn the parameters of the Transformer, where $Y$ is the true label information. Autoregressive language modeling (Radford et al., 2018; Brown et al., 2020) involves using a causal mask $M$, which is a lower triangular matrix, and is added before the softmax operation as follows:
$\operatorname{attn}(X)$ $=\operatorname{softmax}(M+QK^{\top})V,$
$M_{ij}$ $=0\text{ if }i\geq j\text{ else }-\infty$
@@ 3.2 Gradient Analysis of Attention
For simplicity, we first let the sequence length $N$ be $2$ with attention probabilities $A=\operatorname{softmax}(QK^{\top})$ and attention logits as $\tilde{A}=QK^{\top}$. Expanding the matrices $A$ and $\tilde{A}$, we get:
@@
$=\begin{pmatrix}\sigma(\tilde{a}_{11}-\tilde{a}_{12})v_{1}+(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))v_{2}\\ \sigma(\tilde{a}_{21}-\tilde{a}_{22})v_{1}+(1-\sigma(\tilde{a}_{21}-\tilde{a}_{22}))v_{2}\end{pmatrix},$ (4)
@@
Thus, small Jacobian magnitude can lead to small backpropagated gradient. We now analyze an element of $\operatorname{attn}$ Jacobian:
$\frac{\partial o_{1}}{\partial\tilde{a}_{11}}$ $=v_{1}\sigma(\tilde{a}_{11}-\tilde{a}_{12})(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))$
$\quad-v_{2}\sigma(\tilde{a}_{11}-\tilde{a}_{12})(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))$
@@
$\implies\operatorname{laser}(X)$ $=\log(\operatorname{softmax}(QK^{\top})\exp(V))$ (6)
@@
where $\log(.)$ and $\exp(.)$ are applied elementwise. Expanding (6) for $N=2$ and $d=1$ as done for standard attention (4) gives LASER output:
$\alpha_{1}$ $=\sigma(\tilde{a}_{11}-\tilde{a}_{12}),\quad\alpha_{2}=\sigma(\tilde{a}_{21}-\tilde{a}_{22})$
$\begin{pmatrix}o_{1}\\ o_{2}\end{pmatrix}$ $=\begin{pmatrix}\log(\alpha_{1}\exp(v_{1})+(1-\alpha_{1})\exp(v_{2}))\\ \log(\alpha_{2}\exp(v_{1})+(1-\alpha_{2})\exp(v_{2}))\end{pmatrix},$ (7)
Low gradient saturation. Computing an element in the Jacobian ${\partial\operatorname{laser}(X)}/{\partial\tilde{A}}$ as done for standard attention in (5) gives the equation for LASER Jacobian:
$\frac{\partial o_{1}}{\partial\tilde{a}_{11}}$ $=\frac{(\exp(v_{1})-\exp(v_{2}))\alpha_{1}(1-\alpha_{1})}{\alpha_{1}\exp(v_{1})+(1-\alpha_{1})\exp(v_{2})}$
$=\frac{(\exp(v_{1})-\exp(v_{2}))\alpha_{1}(1-\alpha_{1})}{\alpha_{1}(\exp(v_{1})-\exp(v_{2}))+\exp(v_{2})}.$
We now consider a limiting case to simplify the above equation: if $\exp(v_{1})\gg\exp(v_{2})$,
$\frac{\partial o_{1}}{\partial\tilde{a}_{11}}$ $=\frac{\alpha_{1}(1-\alpha_{1})}{\alpha_{1}+\exp(v_{2})/(\exp(v_{1})-\exp(v_{2}))}$
$\approx 1-\alpha_{1}=\underbrace{(1-\sigma(\tilde{a}_{11}-\tilde{a}_{12}))}_{\text{low saturation}},$
@@
Lemma 3.2 ((Boyd & Vandenberghe, 2004)).
@@
$\max\{x_{1},\dots,x_{n}\}\leq f(x_{1},\ldots,x_{n})\leq\max$$\{x_{1},\dots,x_{n}\}$
$+\log n$
@@ 3.3 LASER Implementation via Log-Weighted-Sum-Exp Trick
$\alpha_{1}$ $=\sigma(\tilde{a}_{11}-\tilde{a}_{12}),$
$o_{1}$ $=\log(\alpha_{1}\exp(v_{1})+(1-\alpha_{1})\exp(v_{2})),$
one can notice that $\exp(.)$ operations can lead to overflow. This problem has been recognized in (Blanchard et al., 2019) and the “log-sum-exp trick” is used to avoid overflows. However, the log-sum-exp trick cannot be applied directly as it would be difficult to implement without changing the underlying attention function. We propose a “log-weighted-sum trick”, where we subtract the maximum value $m=\max(v_{1},v_{2})$ from $v_{1}$ and $v_{2}$ and rewrite the above equation as follows:
$o_{1}$ $=\log((\alpha_{1}\exp(v_{1}-m)$
$\quad+(1-\alpha_{1})\exp(v_{2}-m))\cdot\exp(m))$
$=\log(\alpha_{1}\exp(v_{1}-m)+(1-\alpha_{1})\exp(v_{2}-m))+m.$
@@
$O$ $=\log(\operatorname{softmax}(QK^{\top})\exp(\hat{V})\operatorname{diag}(\exp(m)))$
$\quad O_{ij}$ $=(\log(\operatorname{softmax}(QK^{\top})\exp(\hat{V})))_{ij}+m_{j},$ (9)
where $O\in\mathbb{R}^{N\times d}$. Here, $m=(m_{1},\ldots,m_{d})$ and $\operatorname{diag}(m)$ is a diagonal matrix with elements of $m$ as diagonals. Log-weighted-sum-exp trick allows us to implement LASER attention via merely modifying the inputs and outputs of standard attention, without changing the underlying attention function. Additionally, we show in Section 4.1 that this trick helps avoid overflows in large scale settings. The following JAX (Bradbury et al., 2018) code demonstrates that LASER attention can be implemented by utilizing standard attention functions.
@@
// Shift values to avoid overflow in the following
$\hat{V}\in\mathbb{R}^{N\times d}\text{ such that }\hat{V}_{ij}=(V_{ij}-m_{j})$
@@
$O\in\mathbb{R}^{N\times d}$ is such that:
$(O)_{ij}=(\log(\operatorname{softmax}(QK^{\top})\exp(\hat{V})))_{ij}+m_{j}$
6:  Return $O$
@@ 4.1 Autoregressive Language Modeling on C4
Figure 2: Comparison of pretraining performance between LASER and the standard attention mechanism on a 301M parameter autoregressive language model. The model consists of 32 layers, a 2048 hidden dimension, and a 1024 model dimension, trained on 168 billion tokens from the C4 dataset. LASER attention consistently achieves lower training (top) and test (bottom) loss values.
@@
To demonstrate the scalability of our approach, we conducted experiments on a 1.1 billion and 2.2 billion parameter model. We note that without using the log-weighted-sum-exp trick introduced in Section 3.3, the 2.2 billion parameter model training fails due to numerical overflow. In Table 2, we show that LASER attention outperforms standard-attention in a 2.2 billion parameter model with model dimension 2048 and hidden dimension 8192 with 32 layers and 8 attention heads (each of size 512). We also train a 1.1 billion model, which has a scaled down hidden dimension (4096) and attention head size (256). We show the training curves of both the models in Figure 5, Appendix A.3. In Figure 4, we used a power law fit $f(n)=an^{b}$ to fit the final test loss values of LLM training runs as a function of number of parameters. Additionally, we conducted training analysis, by qualitatively measuring stability of our training runs, sensitivity to bfloat16 training in Appendix A.5.
Performance Analysis.
We note that the 2.2B model with standard attention takes 27.22 hours on TPU v5 (Cloud, 2023) to reach a least test loss value of 2.327, however, LASER takes 24.88 hours (relative improvement of 9.4%).
Figure 4: Scaling law: Power-law fit for test loss against number of parameters. This plot uses 234M, 300M, 435M, 1.1B and 2.2B parameter models’ final test losses after training on $\sim$ 167B tokens. To reach a loss of 2.347, it takes 15.65% fewer parameters with LASER attention.
Baselines
Standard dot-product attention formulation in (Vaswani et al., 2017) uses a temperature of $\sqrt{d}$. Lower temperatures would be able to express sharper distributions. Attention function with temperature hyperparameter can be formalized as follows:
Standard+temp: $\quad\operatorname{softmax}\left(\frac{QK^{\top}}{\tau\sqrt{d}}\right)V,$
LASER+temp: $\quad\log\left(\operatorname{softmax}\left(\frac{QK^{\top}}{\tau\sqrt{d}}\right)\exp(V)\right),$
where the temperature hyperparameter $\tau$ is tuned manually. We now introduce per-dim temp operation, which scales each dimension with its own trainable temperature value:
[…]
$D=\operatorname{diag}(\operatorname{softplus}(p))=\operatorname{diag}(\log(1+\exp(p)))\in\mathbb{R}^{d\times d},$
where $p$ is a $d$-dimensional trainable parameter and $\operatorname{diag}(\operatorname{softplus}(\cdot))$ returns a diagonal matrix with positive diagonal entries to be scaled with queries. This technique is implemented in large language modeling frameworks such as PAX (Google, 2025) and AlgoPerf (Dahl et al., 2023), a framework used to compare optimizers.
In large vision transformers (Dosovitskiy et al., 2021), the norms of queries/keys can vary substantially across tokens with sharp softmax distributions in standard attention model training, leading to training instabilities. In (Dehghani et al., 2023), the authors propose QK-Normalization which conducts LayerNorm on queries and keys before computing attention weights:
Standard+QK-Normalization:
$\operatorname{softmax}(\operatorname{LayerNorm}(Q)\operatorname{LayerNorm}(K)^{\top}/\sqrt{d})V$
@@
In Table 2, we mention the performance of our 2.2 billion parameter model on several downstream tasks. Where we evaluate on ARC (Clark et al., 2018), BoolQ (Clark et al., 2019), CB (Wang et al., 2019), COPA (Wang et al., 2019), HellaSwag (Zellers et al., 2019), MultiRC (Khashabi et al., 2018), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), RACE (Lai et al., 2017), ReCoRD (Zhang et al., 2018), RTE (Wang et al., 2019), StoryCloze (Mostafazadeh et al., 2016), WiC (Pilehvar & Camacho-Collados, 2019), Winograd (Levesque et al., 2012), Winogrande (Kocijan et al., 2020), and WSC (Wang et al., 2019). We found that LASER outperforms with upto 3.38% difference and 1% difference on average in accuracy.
(a) Part 1 Dataset +per-dim temp LASER +per-dim temp Standard +temp WSC 81.40 77.89 79.65 76.14 78.25 Winogrande 62.27 61.72 62.35 60.69 62.19 Winograd 81.68 78.75 80.22 80.95 79.85 WiC 52.04 47.02 51.10 47.34 47.18 StoryCloze 77.87 77.77 76.48 76.22 76.17 RTE 53.43 54.51 53.07 54.15 51.70 ReCoRD 85.29 85.05 85.10 84.82 84.89 RaceM 50.42 50.91 49.65 49.09 48.89
(b) Part 2 Dataset +per-dim temp LASER +per-dim temp Standard +temp RaceH 37.88 37.94 37.65 37.62 37.04 PIQA 77.09 77.20 76.88 76.55 76.45 OpenBookQA 48.80 46.80 47.60 47.80 48.48 MultiRC 57.03 57.78 53.94 55.36 53.38 HellaSwag 66.58 66.70 65.42 65.53 65.84 COPA 82.00 87.00 80.00 82.00 85.00 CB 41.07 44.64 42.86 42.86 42.86 BoolQ 63.52 64.43 60.70 62.42 61.69
Table 2: Accuracies of one-shot evaluation of a 2.2 billion parameter autoregressive language model trained via LASER and standard attention. We found that LASER outperforms or performs the same as standard attention by up to 3.4%. On average, LASER gives an accuracy of 63.39% vs standard attention’s 62.34%. We also include a temperature-tuned version of standard attention, which gives an average accuracy of 62.49%. Additionally, we report per-dimension temperature scaling for both LASER and standard attention, achieving overall means of 63.52% and 62.56% respectively.
Finetuning on SuperGLUE.
To further evaluate LASER, we fine-tuned the 2.2B parameter model on the SuperGLUE dataset (Wang et al., 2019) for 10 epochs. These experiments provide a more comprehensive view of how LASER improves over standard attention beyond 1-shot downstream evaluations. The results, presented in Table 3, show a 1.65% average improvement in decoding accuracy for LASER over standard attention.
Table 3: SuperGLUE fine-tuning results for the 2.2B parameter model (decoding accuracy %). LASER shows a 1.65% average improvement over Standard Attention.
Task LASER Standard
COPA 57.00 58.00
WiC 56.64 53.92
WSC 40.38 36.54
RTE 22.02 20.94
Average 44.01 42.35